# Reinforcement Learning from Probabilistic Forecasts for Safe Decision-Making via Conditional Value-at-Risk Planning

### v1

@@ 4   Sample $\theta_{k}\sim b$;
6      Compute $Q(b_{t},s_{t},a)$ under $\theta_{k}$;
@@ 15 $b\leftarrow b_{t+1}$;
16
17
181exTerminate when horizon $(T)$ is reached;
Algorithm 1 UAMDP($b_{0},H,F$)
@@ 3.2.1 Assumption
    Bounded Rewards. Rewards almost surely satisfy $\lvert R_{m}\rvert\leq 1$ and the horizon is $H$.
@@ 3.2.2 Preliminaries
Let $k$ be a fixed episode and condition on $\theta_{k}\sim b_{0}$, and let $M(\theta_{k})$ be the MDP defined as $\langle\mathcal{S},\mathcal{A},P_{\theta},R_{\theta},\gamma\rangle$. Under A3 (Planner Optimality), the algorithm executes the optimal $H$-step policy for the fully observable $M(\theta_{k})$.
@@
Let $\pi^{U}$ be the policy produced by UAMDP. Under A1–A3 (realizability, particle filter with $N$ particles, and depth-$L$ Monte-Carlo tree search), for any $\delta\in(0,1)$, with probability at least $1-\delta$:
$\mathrm{Regret}_{H}\;\leq\;O\!\left(\sqrt{HSA\,\ln\!\frac{1}{\delta}}\;+\;\frac{C_{1}}{\sqrt{N}}\;+\;\frac{C_{2}}{L}\right),$
@@ 3.2.3 Proof of correctness
Before stating the theorem, recall that our algorithm draws a single parameter sample $\theta_{k}\sim b_{0}$ at the start of each episode (A4) and then plans optimally for $H$ using with that sample (A3). Because the true parameter $\theta^{\ast}$ lies within the model class (A1) and filtering is exact (A2), the belief state $(b_{t},s_{t})$ contains all the information needed for optimal control. The next theorem formalises this intuition.
@@
For $(b,s)\in\mathcal{B}\times\mathcal{S}$ and $a\in\mathcal{A}$, define the Bayesian posterior (the belief update) as the map $\mathcal{B}(b,x,a)$ whose density at $\theta$ is
$\mathcal{B}(b,x,a)(\theta)\;:=\;\frac{p_{\phi}(x\mid s,a,\theta)\,b(\theta)}{\int p_{\phi}(x\mid s,a,\vartheta)\,b(\vartheta)\,d\vartheta},$
and define the belief–state kernel by
$\tilde{p}\!\left((b^{\prime},s^{\prime})\mid b,s,a\right)\;=\;\int_{\theta}p_{\phi}(s^{\prime}\mid s,a,\theta)\,\delta_{\mathcal{B}(b,s,a)}(b^{\prime})\,b(\theta)\,d\theta.$
@@
$V_{t}^{\ast}(b,s)\;=\;\sup_{\pi\in\Pi}\;\mathbb{E}\!\left[\sum_{h=t}^{H-1}\gamma^{\,h-t}\,r(s_{h},a_{h})\;\middle|\;b_{t}=b,\;s_{t}=s\right]$
where
$\Pi=\{\pi:\mathcal{B}\times\mathcal{S}\to\mathcal{A}\ \text{measurable}\}.$
For any initial belief–state pair $(b_{0},s_{0})$ and horizon $H$, the expected return achieved by UAMDP, denoted $V^{\mathrm{ALG}}_{0}(b_{0},s_{0})$, coincides with the finite-horizon Bayes value $V^{\ast}_{0}(b_{0},s_{0})$, i.e., $V^{\mathrm{ALG}}_{0}(b_{0},s_{0})=V^{\ast}_{0}(b_{0},s_{0})$. Let $\mathbb{E}_{\theta}$ denote expectation with respect to the initial parameter draw $\theta_{k}\sim b_{0}$. By A3 (planner optimality), for the fixed $\theta_{k}$ the planner selects the unique $H$-step optimal policy of the fully observable MDP
@@
Lemma 1 shows that the sequence $\{(b_{t},s_{t})\}$ is Markov with kernel $\tilde{p}$; hence, every Bayes-optimal policy can be viewed as acting on this Markov chain alone. Combining this with A2 (exact filtering) implies that $V^{\ast}_{0}\!\big(\mathcal{M}(\theta_{k})\big)=V^{\ast}_{0}(b_{0},s_{0})$ almost surely under the draw of $\theta_{k}$. Taking expectation over $\theta_{k}$ and applying the tower property,
@@ 3.2.4 Error bounds
Let $R_{\max}=\sup_{(s,a)}\lvert r(s,a)\rvert$ denote the largest per-step reward. We quantify the two inaccuracies by scalars $\varepsilon_{f},\varepsilon_{p}\in\bigl[0,\tfrac{1}{2}R_{m}\bigr]$, such that $\varepsilon_{f}$ denote the maximum single-step filter and $\varepsilon_{p}$ is the planner errors. Since the chosen action $a_{t}$ satisfies
$\bigl\lvert Q^{*}_{\theta,\,H-t}(s_{t},a_{t})-\max_{a}Q^{*}_{\theta,\,H-t}(s_{t},a)\bigr\rvert\leq\varepsilon_{p},$
its discounted contribution to the value gap is at most $\gamma^{t}\,\varepsilon_{p}$. By belief Lipschitz lemma [39], for any two beliefs $b,b^{\prime}$ and state $s$, it follows that
$\bigl|V_{t}^{*}(b,s)-V_{t}^{*}(b^{\prime},s)\bigr|\leq\frac{2\gamma^{H-t}R_{m}}{1-\gamma}\,\|b-b^{\prime}\|_{1}.$
This implies that replacing the true belief $b_{t}$ with the approximation $\hat{b}_{t}$ changes the optimal value by at most $\frac{2\gamma^{H-t}R_{m}\varepsilon_{f}}{1-\gamma}$, so the discounted impact on step $t$ is $\gamma^{t}\!\left(\frac{2\gamma^{H-t}R_{m}\varepsilon_{f}}{1-\gamma}\right)$. Adding the two sources and summing the resulting geometric series gives:
$\Delta_{0}$ $=V_{0}^{*}(b_{0},s_{0})-V_{0}^{\mathrm{ALG}}(b_{0},s_{0})$
$\leq\sum_{t=0}^{\infty}\gamma^{t}\varepsilon_{p}+\sum_{t=0}^{\infty}\gamma^{t}\frac{2\gamma^{H-t}R_{m}\varepsilon_{f}}{1-\gamma}$
$=\frac{\varepsilon_{p}}{1-\gamma}+\frac{2\gamma R_{m}\varepsilon_{f}}{(1-\gamma)^{2}}=\frac{\gamma}{1-\gamma}\left(\varepsilon_{p}+\frac{2\gamma R_{m}\varepsilon_{f}}{1-\gamma}\right).$
Given the mild condition is holds in all the experimental settings,
$2\gamma R_{m}\leq 1-\gamma,$
The error is bounded by:
$\Delta_{0}\leq\frac{\gamma}{1-\gamma}\left(\varepsilon_{p}+\varepsilon_{f}\right).$
@@ 4.3 Experimental procedure
  • Preprocessing and feature construction. As detailed in Section 4.2, domain-specific signals were extracted and normalized to ensure consistency across time-series models.
@@
  • Model training and selection. Six probabilistic forecasters (ARIMA, DeepAR, GPs, TFT, PatchTST, and Prophet) were trained on the training set and early-stopped based on validation using the continuous ranked probability score (CRPS). Hyperparameters for each method were selected via Bayesian optimization with a budget of 50 trials.
  • Risk-aware reinforcement learning setup.
    • For the S&P 500 dataset, we designed a portfolio management gym with three discrete actions {cash, long index, bond ETF}, a transaction cost of $0.02\%$, and daily rebalancing.
    • For the H&M dataset, we implemented an inventory control gym, where actions correspond to reorder quantities and rewards incorporate margin, holding cost, and stock-out penalties.
@@
    • We assessed hyperparameter stability by evaluating the agent across $\mathrm{CVaR}$ confidence levels $\alpha\in\{0.01,0.05,0.10\}$ and risk weights $\lambda$ $\in\{0$, $0.3$, $0.6$, $0.8$, 1.0}1.0\}.
@@ 4.4 Demonstration
The portfolio starts at \$1,000, composed of 50% cash and 50% index. Transaction costs are $k=0.02\%$ per trade. The forecaster is a GP with observation noise $\sigma_{\varepsilon}^{2}=2.5\times 10^{-4}$. The $\mathrm{CVaR}$ head uses confidence level $\alpha=0.05$ and risk weight $\lambda=0.7$. At $t=0$, the GP predicts next-period log-return $r_{1}\sim\mathcal{N}(\mu_{0}=0,\ \sigma_{0}^{2}=5\times 10^{-4})$. UAMDP draws a scenario $\tilde{r}_{1}=0.009$ ($\approx 0.9\%$ up-move) and feeds it to the planner. Solving the $\mathrm{CVaR}$ objective yields a switch to 80% equity, 20% cash. The realized return is $r_{1}=0.0198$. The GP posterior becomes:

### v2

@@ 4   Sample $\theta_{k}\sim b$;
6    if $n\geq T$ then
7       break
8     Estimate $Q(b_{t},s_{t},a)$ under $\theta_{k}$;
@@ 18 $b\leftarrow b_{t+1}$;
19
Algorithm 1 UAMDP($b_{0},H,T,F$): online control with Bayesian belief update.
such that parameter values explaining the data gain posterior mass, and future decisions become correspondingly sharper.
Let $(\Theta,\mathcal{T})$ be the measurable space of latent environment parameters and let $\lambda$ be a $\sigma$-finite reference measure on $\Theta$ (Lebesgue for continuous $\Theta$, counting measure for discrete $\Theta$). We represent beliefs by densities w.r.t. $\lambda$ and define the belief space
$\mathcal{B}:=\left\{b:\Theta\to\mathbb{R}_{\geq 0}\ \middle|\ \int_{\Theta}b(\theta)\,\lambda(d\theta)=1\right\}.$
Equivalently, each $b\in\mathcal{B}$ induces a probability measure $b(d\theta)=b(\theta)\lambda(d\theta)$. Throughout, $p_{\phi}(x\mid s,a,\theta)$ denotes the forecaster’s predictive density and is assumed measurable in $\theta$ and integrable so that the normalizing constant below is finite and non-zero. For $(b,s)\in\mathcal{B}\times\mathcal{S}$, $a\in\mathcal{A}$, and observation $x\in\mathbb{R}^{d}$, we define the normalized Bayesian update operator $B:\mathcal{B}\times\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{d}\to\mathcal{B}$ by
$B(b,s,a,x)(\theta):=\frac{p_{\phi}(x\mid s,a,\theta)\,b(\theta)}{\int_{\Theta}p_{\phi}(x\mid s,a,\vartheta)\,b(\vartheta)\,\lambda(d\vartheta)}.$
Accordingly, the proportional update above is $b_{t+1}=B(b_{t},s_{t},a_{t}^{\star},x_{t+1})$.
Optionally, a risk-sensitive layer is interposed before action execution. In that case, the planner replaces the mean-based utility with a coherent risk measure. Let
$Z_{t}\;=\;\sum_{h=0}^{H-1}\gamma^{h}\,r_{t+h}$
denote the random $H$-step discounted return generated from $(b_{t},s_{t},a)$ under the current policy and dynamics. The distributional critic $Z_{\psi}(b_{t},s_{t},a)$ parameterizes an approximation to the law of $Z_{t}$, and we evaluate $\mathrm{CVaR}_{\alpha}[Z_{\psi}(b_{t},s_{t},a)]$ in lieu of $Q(b_{t},s_{t},a)$. Here $\mathrm{CVaR}_{\alpha}$ is the conditional value-at-risk at level $\alpha\in(0,1)$, defined as the average of the worst $\alpha$-fraction of outcomes in the return distribution. In implementation, this quantity is computed via the standard empirical estimator: given a batch of Monte-Carlo samples from $Z_{\psi}(b_{t},s_{t},a)$, we sort the samples and average those in the lower $\alpha$-tail. Within both MPC and MCTS, the planning logic is unchanged; only the leaf-evaluation or trajectory objective is modified so that action sequences are ranked according to their CVaR-adjusted $H$-step returns rather than their expected returns. Alternatively, when a domain-specific safety set $S_{\mathrm{safe}}\subseteq\mathbb{R}^{d}$ and tolerance $\delta$ are provided, the risk layer enforces chance constraints
thereby guaranteeing probabilistic safety at the constraint level.
Finally, the loop of sampling, planning, execution, and Bayesian updating repeats until the fixed interaction horizon $T$ is reached or an earlier task-specific termination criterion is met. Through this closed-loop interaction, the UAMDP continuously couples uncertainty-aware forecasting with decision-making, ensuring that every action both exploits current knowledge and refines future understanding.
@@ 3.2.1 Assumption
    Bounded Rewards. For all $(s,a)$ we have $\lvert r(s,a)\rvert\leq 1$ and the horizon is $H$.
@@ 3.2.2 Preliminaries
Let $k$ be a fixed episode and condition on $\theta_{k}\sim b_{0}$, and let $M(\theta_{k})$ be the MDP defined as $\langle\mathcal{S},\mathcal{A},P_{\theta_{k}},r,\gamma\rangle$. Under A3 (Planner Optimality), the algorithm executes the optimal $H$-step policy for the fully observable $M(\theta_{k})$.
@@
Let $\pi^{U}$ be the policy produced by UAMDP. Under A1–A3 (realizability, a particle filter with $N$ particles, and a depth-$L$ Monte Carlo tree search planner), for any $\delta\in(0,1)$, with probability at least $1-\delta$:
$\mathrm{regret}_{H}\;\leq\;\mathcal{O}\!\left(\sqrt{H\,|\mathcal{S}|\,|\mathcal{A}|\,\ln\!\frac{1}{\delta}}\;+\;\frac{C_{1}}{\sqrt{N}}\;+\;\frac{C_{2}}{L}\right),$
@@
It is worth noting that the regret bound in Theorem 2 assumes a planner that is bounded to depth $L$ at each decision epoch. In our empirical study, the MCTS planner described in Section 4.3 uses a fixed finite search depth, so this assumption is satisfied and the bound qualitatively captures how regret scales with the planning horizon $H$, the number of particles $N$, and the search depth $L$. We do not treat the bound as a tight numerical prediction but rather as a guide to the effect of these approximation parameters.
@@ 3.2.3 Proof of correctness
Before stating the theorem, recall that our algorithm draws a single parameter sample $\theta_{k}\sim b_{0}$ at the start of each episode (A4) and then plans optimally for $H$ using that sample (A3). Because the true parameter $\theta^{\ast}$ lies within the model class (A1) and filtering is exact (A2), the belief state $(b_{t},s_{t})$ contains all the information needed for optimal control. The next theorem formalises this intuition.
@@
Recall the belief space $\mathcal{B}$ and Bayesian update operator $B$ defined above. Let $\delta_{y}$ denote the Dirac measure concentrated at $y$. We then define the belief–state kernel by
$\tilde{p}\!\left((b^{\prime},s^{\prime})\mid b,s,a\right)$ $=\int_{\Theta}\!\int_{\mathbb{R}^{d}}p_{\phi}(x\mid s,a,\theta)\,\delta_{B(b,s,a,x)}(b^{\prime})\,\delta_{\tau(s,a,x)}(s^{\prime})\,b(\theta)\,\lambda(d\theta)\,dx.$
Let $\Delta(\mathcal{A})$ denote the set of probability measures on $\mathcal{A}$ and let
$\Pi:=\left\{\pi:\mathcal{B}\times\mathcal{S}\to\Delta(\mathcal{A})\ \text{measurable}\right\}.$
For $\pi\in\Pi$, actions are sampled as $a_{h}\sim\pi(\cdot\mid b_{h},s_{h})$. Deterministic policies are included as a special case via $\pi(\cdot\mid b,s)=\delta_{\bar{a}(b,s)}(\cdot)$.
@@
$V_{t}^{\ast}(b,s)\;=\;\sup_{\pi\in\Pi}\;\mathbb{E}^{\pi}\!\left[\sum_{h=t}^{H-1}\gamma^{\,h-t}\,r(s_{h},a_{h})\;\middle|\;b_{t}=b,\;s_{t}=s\right].$
For any initial belief–state pair $(b_{0},s_{0})$ and horizon $H$, the expected return achieved by UAMDP, denoted $V^{\mathrm{ALG}}_{0}(b_{0},s_{0})$, coincides with the finite-horizon Bayes value $V^{\ast}_{0}(b_{0},s_{0})$, i.e., $V^{\mathrm{ALG}}_{0}(b_{0},s_{0})=V^{\ast}_{0}(b_{0},s_{0})$. Let $\mathbb{E}_{\theta}$ denote expectation with respect to the initial parameter draw $\theta_{k}\sim b_{0}$. By A3 (planner optimality), for the fixed $\theta_{k}$ the planner selects the $H$-step optimal policy of the fully observable MDP
@@
Lemma 1 shows that the sequence $\{(b_{t},s_{t})\}$ is Markov with kernel $\tilde{p}$; hence every Bayes-optimal policy can be viewed as acting on this Markov chain alone. Combining this with A2 (exact filtering) and the definition of the Bayes value yields
@@ 3.2.4 Error bounds
Let $R_{\max}=\sup_{s,a}|r(s,a)|$ denote the largest per-step reward. We quantify the two inaccuracies by scalars $\varepsilon_{f},\varepsilon_{p}\geq 0$, where $\varepsilon_{f}$ bounds the maximum single-step filter error and $\varepsilon_{p}$ bounds the planner error. Since the chosen action $a_{t}$ satisfies
$\bigl|Q^{\star}_{\theta,H-t}(s_{t},a_{t})-\max_{a}Q^{\star}_{\theta,H-t}(s_{t},a)\bigr|\leq\varepsilon_{p},$
its discounted contribution to the value gap is at most $\gamma^{t}\varepsilon_{p}$. By the belief Lipschitz lemma [russo2014learning], for any two beliefs $b,b^{\prime}$ and state $s$, it follows that
$\bigl|V^{\star}_{t}(b,s)-V^{\star}_{t}(b^{\prime},s)\bigr|\;\leq\;\frac{2\gamma^{H-t}R_{\max}}{1-\gamma}\,\|b-b^{\prime}\|_{1}.$
This implies that replacing the true belief $b_{t}$ with the approximation $\hat{b}_{t}$ changes the optimal value by at most $\frac{2\gamma^{H-t}R_{\max}}{1-\gamma}\,\varepsilon_{f}$, so the discounted impact at step $t$ is
$\gamma^{t}\,\frac{2\gamma^{H-t}R_{\max}}{1-\gamma}\,\varepsilon_{f}.$
Adding the two sources and summing the resulting geometric series gives
$\Delta_{0}$ $=V^{\ast}_{0}(b_{0},s_{0})-V^{\mathrm{ALG}}_{0}(b_{0},s_{0})$
$\leq\sum_{t=0}^{\infty}\gamma^{t}\,\varepsilon_{p}+\sum_{t=0}^{\infty}\gamma^{t}\,\frac{2\gamma^{H-t}R_{\max}}{1-\gamma}\,\varepsilon_{f}$
$=\frac{\varepsilon_{p}}{1-\gamma}+\frac{2\gamma R_{\max}}{(1-\gamma)^{2}}\,\varepsilon_{f}.$ (1)
Thus, for arbitrary $\gamma\in(0,1)$ and finite $R_{\max}$, the gap between the ideal Bayes value and the value achieved with approximate filtering and planning is bounded by
$\Delta_{0}\;\leq\;\frac{\varepsilon_{p}}{1-\gamma}+\frac{2\gamma R_{\max}}{(1-\gamma)^{2}}\,\varepsilon_{f}.$
In particular, when rewards are normalized so that $R_{\max}=1$ and $\gamma$ is close to one, the leading dependence is of order $\mathcal{O}\!\bigl((\varepsilon_{p}+\varepsilon_{f})/(1-\gamma)^{2}\bigr)$, reflecting the fact that small per-step approximation errors can accumulate over long effective horizons.
@@ 4.2 Pre-processing and feature engineering
To ensure fairness across methods, the same feature vectors constructed above are supplied to all probabilistic forecasters and reinforcement learning agents, including the baselines. The UAMDP framework itself is agnostic to the particular choice of hand-crafted signals: it only requires that the forecaster consumes a covariate vector $x_{t}$ and returns a calibrated predictive distribution. The S&P 500 and H&M feature sets described in this section therefore reflect standard practice in their respective domains rather than a requirement of the framework, and any performance differences between UAMDP and the baselines arise from the forecasting and control components rather than from privileged access to domain-specific features.
Although the engineered feature sets differ between the S&P 500 and H&M domains, they are applied uniformly within each domain. For a given dataset, every probabilistic forecasting model uses the same normalized feature vector $x_{t}$, and the associated RL agents operate on the same augmented state $s_{t}=(x_{1:t},a_{0:t-1})$ defined in Table 1. Consequently, performance differences between UAMDP and the baselines reflect how each method models and exploits predictive distributions for control, rather than differences in feature engineering.
@@ 4.3 Experimental procedure
  • Pre-processing and feature construction. As detailed in Section 4.2, domain-specific signals were extracted and normalized to ensure consistency across time-series models.
@@
  • Model training and selection. Six probabilistic forecasters (ARIMA, DeepAR, GPs, TFT, PatchTST, and Prophet) were trained independently of the RL agent on the training set by minimizing the continuous ranked probability score (CRPS) and early-stopped based on validation CRPS. Hyperparameters for each method were selected via Bayesian optimization with a budget of 50 trials. After training, the forecaster parameters $\phi$ were frozen and used as fixed likelihood modules during subsequent RL training and evaluation on the temporally held-out test set described above.
  • Risk-aware reinforcement learning setup. In both environments we instantiate the abstract state $s_{t}$ from Table 1 as the augmented history $s_{t}=(x_{1:t},a_{0:t-1})$, where $x_{t}$ is the domain-specific feature vector from Section 4.2 and $a_{t}$ is the action taken at time $t$. Concretely:
    • S&P 500 (trading). At each trading day $t$ the agent observes the feature vector $x_{t}$ containing the price-, volatility-, volume-, and calendar-based signals defined in Section 4.2, together with the past actions encoded in $s_{t}$. It then selects an action $a_{t}\in\{\text{cash},\text{long index},\text{bond ETF}\}$, which specifies the portfolio allocation for the next period. The one-step reward $r_{t}=r(s_{t},a_{t})$ is the realized daily log-return of the portfolio, as defined in Section 4.2, net of a transaction cost of $0.02\%$ applied to changes in position.
    • H&M (inventory). At each decision epoch the agent observes the demand and price covariates $x_{t}$ described in Section 4.2 (growth of demand, relative price changes, promotion flags, seasonality and lagged indicators). The action $a_{t}$ is the reorder quantity issued to the inventory simulator. The per-period reward $r_{t}=r(s_{t},a_{t})$ aggregates sales margin with holding costs for excess inventory and stock-out penalties whenever demand exceeds available supply, as specified in the inventory control gym.
  • Planning backend and CVaR integration. In all experiments, the planner in Fig. 1 and Algorithm 1 is realized as a bounded-depth Monte Carlo tree search (MCTS); the MPC backend is included for completeness but is not used in the reported results. At each decision step, MCTS expands a tree up to a fixed maximum depth $L$ and evaluates leaf nodes using the empirical $\mathrm{CVaR}_{\alpha}$ objective defined in Section 3.1: given Monte Carlo samples from the distributional critic $Z_{\psi}(b_{t},s_{t},a)$, we compute the empirical $\mathrm{CVaR}_{\alpha}$ and use this value as the leaf estimate. Standard MCTS backup rules then propagate these CVaR-adjusted values to the root to rank candidate actions. In both the S&P 500 and H&M environments, we keep the chance-constraint primitive inactive ($S_{\mathrm{safe}}=\emptyset$), encoding transaction costs and stock-out penalties directly in the reward function and environment dynamics.
@@
    • We assessed hyperparameter stability by evaluating the agent across $\mathrm{CVaR}$ confidence levels $\alpha\in\{0.01,0.05,0.10\}$ and risk weights $\eta\in\{0,0.3,0.6,0.8,1.0\}$.
  • Evaluation metrics and reporting. For the S&P 500 trading environment, performance is evaluated on the sequence of daily portfolio returns induced by the agent’s actions. We report the mean daily Sharpe ratio, maximum drawdown, turnover, and the fraction of days with positive return, as shown in Table 11. The daily returns used in these metrics are based on the log-returns $r_{t}$ defined in Section 4.2. Turnover is reported as the average per-day portfolio rebalancing volume, expressed as the fraction of capital traded at each rebalancing step. For the H&M inventory environment, we report service level, stock-out frequency, and gross margin return on investment (GMROI) aggregated over the full test horizon, as discussed in Section 5. Sharpe ratios are computed on the sequence of daily log-returns and are reported on the daily scale. All metrics are evaluated on the fixed temporal splits described above, and the combination of dataset partitions, hyperparameter search budgets, and environment definitions in Sections 4.2–4.3 provides a fully specified protocol for reproducing the reported results.
@@ 4.4 Demonstration
The portfolio starts at \$1,000, composed of 50% cash and 50% index. Transaction costs are $k=0.02\%$ per trade. The forecaster is a GP with observation noise $\sigma_{\varepsilon}^{2}=2.5\times 10^{-4}$. The $\mathrm{CVaR}$ head uses confidence level $\alpha=0.05$ and risk weight $\eta=0.7$. At $t=0$, the GP predicts next-period log-return $r_{1}\sim\mathcal{N}(\mu_{0}=0,\ \sigma_{0}^{2}=5\times 10^{-4})$. UAMDP draws a scenario $\tilde{r}_{1}=0.009$ ($\approx 0.9\%$ up-move) and feeds it to the planner. Solving the $\mathrm{CVaR}$ objective yields a switch to 80% equity, 20% cash. The realized return is $r_{1}=0.0198$. The GP posterior becomes: